OpenStudy (anonymous):

find the area of a regular polygon with s sides of length 6 and apothem a

OpenStudy (jhannybean):

The area of a polygon is defined by the equation A=(AP)/2 where apothem = a , perimeter = 6s A=(a(6a))/2 Are you able to simplify this?

OpenStudy (anonymous):

yes thank you!

OpenStudy (jhannybean):

Sweet